Water at 4 degrees Celsius has a density of 1000 kg/m^3. What is the equivalent value in units of:

a. lbm/ft^3
b. gmol/L

Please show all work and conversions.
Answer #1

Answer(a)

Density of water ρ =1000 kg/m3 (at 4oC) to g/L and lb/ft3

ρ = {1000 (kg /m3)} x {1000 gm / 1 kg } x {1 m / 3.2808}3

= 62.43 lb/ft3

Answer (b)

ρ = {1000 (kg /m3)} x {2.2046 lb / 1 kg } x {1 m3 / 1000 L } = 1000 g/L = (1000 g /L ) / (18 g/mol) = 55.56 mol /L